# Efficient Input-Constrained Impulsive Optimal Control of Linear Systems with Application to Spacecraft Relative Motion

### v1

@@ III-B Input Reconstruction
The technique used to recover the optimal input to the unconstrained problem from the dual formulation can be similarly applied to its constrained version. If $\bm{\lambda}_{f}^{*},\bm{\sigma}^{*}$ are the optimal solution variables to Problem 13, then the optimal impulse times for a given window are recovered from the set
@@
Furthermore, from the duality formulation, it must hold that
$\sigma_{k}\left(\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t-\Delta V_{max,k}\right)=0\ \forall k\in\mathbb{Z}_{1}^{N},$ (17)
@@
$\sigma_{k}>0$ $\iff\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t=\Delta V_{max,k}$ (18)
$\iff\sum_{j=1}^{m_{k}}f_{k}(\bm{v}_{kj})=\Delta V_{max,k},$
@@
From the optimal impulse times, the optimal input can be recovered from the problem:
@@
Solve the discretized primal Problem 10 from the optimal impulse times recovered from Problem 13 with $\mathcal{P}_{d}=$
$\min_{\{\bm{v}\}}$ $\sum_{k=1}^{N}\sum_{j=1}^{m_{k}}f_{k}(\bm{v}_{kj})$ (19)
s.t. $\footnotesize{\bm{x}_{f}=\Phi(t_{f},t_{0})\bm{x}_{0}+\sum_{k=1}^{N}\sum_{j=1}^{m_{k}}\Phi(t_{f},t_{kj})B(t_{kj})\bm{v}_{kj},}$
$\sum_{j=1}^{m_{k}}f_{k}(\bm{v}_{kj})\leq\Delta V_{max,k}\qquad\forall k\in\mathbb{Z}_{1}^{N}.$
@@ IV Algorithmic Approach
Solving Problem 10 in an algorithmically tractable manner can be done through first solving the dual problem to determine the optimal impulse locations and then reconstructing the input through the discretized primal. However, certain nuances must be considered to solve these problems efficiently and robustly.
@@ IV-A Dual Problem Algorithm
Problem 13 in its native form is not readily solvable due to the semi-infinite constraint. Several techniques exist to solve semi-infinite programs (SIPs), but this work and typical applications to spacecraft rendezvous leverage discretization methods, which convert the problem to an iterative finite optimization problem. Moreover, in contrast to Problem 7, which has no possibility of being unbounded above, Problem 13 can be unbounded, corresponding to primal infeasibility. When applying the iterative discretization technique, it is therefore necessary to ensure that premature unboundedness does not produce artificial infeasibility of the problem.
To illustrate the iterative technique to solving the semi-infinite problem, consider the finite form of Problem 13, given by $\mathcal{D}(S)=$
$\max_{\bm{\lambda}_{f},\bm{\sigma}\geq 0}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\sigma_{k}$ (20)
$\forall t\in S\cap T_{k},\ \forall k\in\mathbb{Z}_{1}^{N}$
$\|\bm{\lambda}_{f}\|_{2}\leq\lambda_{max}$
where $S\triangleq\{s_{1},s_{2},\dots s_{d}\}\subset\bigcup_{k=1}^{N}T_{k}$ is a finite set. Since Problem 13 is reducible according to Theorem III.2, then it holds that the finite version of the problem can solve the semi-infinite problem. To leverage this property, discretization techniques gradually add values to the set $S$ until $\mathcal{D}(S)\to\mathcal{D}$. This can be done recursively at each solution iteration, given the optimal solution of the previous iteration $\bm{\lambda}_{f}^{(j)},\bm{\sigma}^{(j)},S^{(j)}$ with
$S^{(j+1)}\leftarrow S^{(j)}\cup\bigcup_{k=1}^{N}\arg\max_{t\in T_{k}}f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{(j)}).$ (21)
The discrete set can be initialized simply with $S^{(1)}=\bigcup_{k=1}^{N}\partial T_{k}$ or with the optimal times associated with a two-impulse burn as is done in [18].
The finite optimization problem in Equation 20 also includes a norm constraint $\|\bm{\lambda}_{f}\|_{2}\leq\lambda_{max}$ to protect against artificial infeasibility in the case that the discretized problem is unbounded. Upon algorithmic convergence, the activeness of this constraint determines whether the semi-infinite problem is unbounded, in which case the primal is infeasible.
These algorithmic modifications retain several important features of discretization methods. Namely, it is evident that $\mathcal{D}\leq\mathcal{D}(S^{(j+1)})\leq\mathcal{D}(S^{(j)})$ since $S^{(j)}\subseteq S^{(j+1)}$. Moreover, given a tolerance $\epsilon$, an important result is as follows, which is extended from [18].
@@
Suppose Equation 20 is solved such that $f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})-\sigma_{k}^{*}\leq 1+\epsilon\ \forall t\in T_{k},\forall k\in\mathbb{Z}_{1}^{N}$. Then it follows that $\mathcal{D}(S^{(j)})\leq(1+\epsilon)\mathcal{D}$.
@@
Let $\bar{\bm{\lambda}}_{f}=\bm{\lambda}_{f}^{*}/(1+\epsilon)$ and $\bar{\sigma}_{k}=\sigma_{k}^{*}/(1+\epsilon)$. Evidently, $(\bar{\bm{\lambda}}_{f},\bar{\bm{\sigma}})$ is a feasible point to Problem 13 that satisfies $\mathcal{D}(S^{(j)})/(1+\epsilon)=\bar{\bm{\lambda}}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\bar{\sigma}_{k}\leq\mathcal{D}\implies\mathcal{D}(S^{(j)})\leq(1+\epsilon)\mathcal{D}$. ∎
@@
1: Discretize $\{T_{k}\}_{k=1}^{N}$
[…]
3:while $\exists k\in\mathbb{Z}_{1}^{N},t\in T_{k}\ \text{s.t.}\ f^{\circ}_{k}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})-\sigma_{k}^{*}>1+\epsilon$ do
4:  $\bm{\lambda}_{f}^{(j)},\bm{\sigma}^{(j)},\mathcal{D}(S^{(j)})\leftarrowSolve$ Optimization Problem 20 with $S^{(j)}$
5:  $S^{(j+1)}\leftarrow S^{(j)}\cup\bigcup_{k=1}^{N}\arg\max_{t\in T_{k}}f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{(j)})$
6:  $j\leftarrow j+1$
7:end while
8:if $\|\bm{\lambda}_{f}^{(j)}\|_{2}=\lambda_{max}$ then
9:  return Problem Infeasible
10:end if
@@ V Performance and Validation
The problem formulation and corresponding algorithmic approach is validated on the highly relevant problem of impulsive control for proximity operations of satellites.
@@ V-A System Definition
Relative motion of satellites in orbits subject to single body gravitational force can be modeled as linear time-varying systems under a variety of additional perturbations, such as drag or spherical harmonics, through a relative orbital element state representation [24, 25], defined by
@@
These states are functions of the classical Keplerian orbital elements of circular two-body motion of the target and chaser spacecraft. Notably, they represent relative motion in terms of integration constants and convert the system to its Jordan canonical form. The linearized equations of motion under this state representation are given by
[…]
Here, $n\triangleq\sqrt{\frac{\mu}{a^{3}}}$ is the orbital mean motion defined based on the planetary gravitational constant ($\mu=398,600.4418\text{km}^{3}/\text{s}^{2}$ for Earth) and the semi-major axis of the orbit ($a=[6538,8378]$ km for LEO). Additionally, $u_{0}$ is the initial mean argument of latitude, which is set to 0 for all ensuing examples. Under these equations of motion, the state transition matrix is simply
$\Phi(t_{f},t_{0})=A(t_{f}-t_{0})+I_{6\times 6}$ (24)
Finally, relative orbital elements can be linearly transformed to relative position and relative velocities in the radial-transverse-normal frame of the target spacecraft using the similarity transformation given by $\delta\bm{x}(t)\triangleq[x,y,z,v_{x},v_{y},v_{z}]^{\intercal}=R(t)\delta\textit{{\oe }}(t)$ where
$R(t)=\begin{bmatrix}1&0&-\text{c}(q)&-\text{s}(q)&0&0\\ 0&1&2\text{s}(q)&-2\text{c}(q)&0&0\\ 0&0&0&0&\text{s}(q)&-\text{c}(q)\\ 0&0&\text{s}(q)&-\text{c}(q)&0&0\\ -\frac{3n}{2}&0&2\text{c}(q)&2\text{s}(q)&0&0\\ 0&0&0&0&\text{c}(q)&\text{s}(q)\end{bmatrix}$ (25)
and $q=nt+u_{0}$.
@@ V-B Planar Reconfiguration Demonstration
The first example considers a planar rendezvous example consisting of reducing along track separation from a distance of 100 km to 10 km. The initial and final conditions are given in Table I. The transfer occurs over four orbits, with eight equally spaced time windows enforced over the duration of the transfer with a buffer of $1/10$th of an orbit between each window. Each time window is discretized over 1000 time points in the dual optimization scheme to provide sufficient resolution. The input is scaled such that a unitary impulse corresponds to $1.13$ [m/s] of $\Delta V$ expenditure. Moreover, the norm applied is the $L_{2}$ norm, such that $f_{k}(\cdot)=\|\cdot\|_{2}\ \forall k$. Finally, the orbital mean motion is $n=.00113$.
TABLE I: Initial Conditions for Planar Reconfiguration (km)
$a\delta a$ $a\delta\lambda$ $a\delta e_{x}$ $a\delta e_{y}$ $a\delta i_{x}$ $a\delta i_{y}$
$\delta\textit{{\oe }}_{0}$ 10 100 0 10 0 0
$\delta\textit{{\oe }}_{f}$ 0 10 0 0 0 0
(a) Relative Position
(b) Throttle
(c) Dual Norm
(d) $\bm{\lambda}_{v}(t)$
Figure 1: Planar Reconfiguration Demonstration.
(a) $\delta a-\delta\lambda$ Trajectory
(b) $\delta e_{x}-\delta e_{y}$ Trajectory
@@
Figure 1 displays the trajectory corresponding to the computed planar reconfiguration control signal and Figure 2 shows the trajectory from the perspective of the relative orbital elements. The computed dual optimal solution for the problem is $\bm{\lambda}_{f}^{*}=[0.7380,0.0891,-0.2359,-0.7236,0,0]^{\intercal},\bm{\sigma}^{*}=[5.4503,3.0092,3.7714,1.3608,2.0926,0,0.4142,1.5183]^{\intercal}$. The total $\Delta V$ expenditure is $\Delta V_{max}\sum_{k=1}^{N}\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t=8.817$ [m/s]. Figure 1a shows the relative positions on the radial-transverse plane with impulses occurring at the blue dots. Figure 1b displays the throttle profile of the computed input signal in blue and time windows shown in red. Evidently, the time windows where the input signal is saturated, i.e. $\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t=1$, correspond to when $\sigma_{k}$ is non-zero. Figure 1c displays the dual norm associated with the semi-infinite constraint and the times at which the dual norm is maximized for each time window. The dual norm exceeds the value of 1 at all but one of the time windows, corresponding exactly to the time windows at which the input signal is saturated. Finally, Figure 1d plots the trajectory formed by $B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}$, which equates to the primer vector $\bm{\lambda}_{v}(t)$ [14], and corresponding impulse locations in blue. Additionally, the unit circle is plotted, corresponding to the sublevel set of the semi-infinite constraint at $\sigma=0$.
@@
To validate and compare the performance of the proposed algorithm, the problem is also solved under a direct optimization framework whereby the primal problem is discretized with the inputs at each discrete node as solution variables. Under this approach, the number of solution variables drastically increases compared to the dual solution framework, with the total number of solution variables amounting to $n_{u}NM$, with $n_{u}$ being the input dimension, $N$ the number of time windows, and $M$ the discretization count per time window. In contrast, the dual SIP involves $n_{x}+N$ solution variables, with cone count being on the order of $\sim 2N$-$3N$, depending on the problem structure. The poor scalability of direct primal optimization manifests as longer solver times, with this issue being especially apparent for long duration transfers, whereby the discretization induces substantial variable counts. This issue is especially prevalent for real-time optimal control of CubeSats, for which CPU performance and memory are limited.
@@
Under this direct optimization approach, the equivalent problem is solved and produces an equivalent solution. However, the solver clock time is 0.6 seconds, whereas the cumulative solver time for the dual SIP with input reconstruction is 0.1262 seconds. Runtime experiments were performed on an Intel i7‑10750H CPU in MATLAB R2024a with the MOSEK solver [26] and YALMIP modeler [27]. Solver times are determined from extracting the MOSEK solver time from the YALMIP solution. Runtime disparities between the proposed approach and direct primal optimization are likely to be more severe for longer duration transfers, where node counts are large, and less severe for shorter duration transfers.
V-C VISORS Transfer Demonstration

### v2

@@ III-B Input Reconstruction
The technique used to recover the optimal input to the unconstrained problem from the dual formulation can be similarly applied to its constrained version. If $\bm{\lambda}_{f}^{*},\bm{\sigma}^{*}$ are the optimal solution variables to Problem 10, then the optimal impulse times for a given window are recovered from the set
@@
Furthermore, from the complementary slackness condition, it must hold that
$\sigma_{k}\left(\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t-\Delta V_{k}\right)=0\ \forall k\in\mathbb{Z}_{1}^{N},$ (13)
@@
$\sigma_{k}>0\implies\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t=\sum_{j=1}^{m_{k}}f_{k}(\bm{v}_{kj})=\Delta V_{k}$ (14)
@@
From the optimal impulse times, the optimal input can be recovered from the tractable finite LP/SOCP:
@@
Solve the discretized primal Problem 4 from the optimal impulse times recovered from Problem 10 with $\mathcal{P}_{d}=\mathcal{P}(\{\{t\}_{k}^{*}\}_{k=1}^{N})$.
Theorem III.1
A feasible optimal solution to Problem 4 consists of at most $n_{x}+N$ impulses.
Proof:
Since strong duality holds and the dual form is a convex semi-infinite program of $n_{x}+N$ variables, Theorem 3.2 of [22] states that there exists a discretization $S$ which satisfies $|S|\leq n_{x}+N$ such that $\mathcal{D}(S)=\mathcal{D}$, leading to at most $n_{x}+N$ points or atoms in time at which $\bm{u}(t)$ may be nonzero. ∎
Remark 1
The number of impulses can be further reduced when considering the addition of the $\bm{\sigma}\geq 0$ constraints, which reduces the active support of the semi-infinite constraint, leading to at most $n_{x}+N_{a}$ points in time at which the input may be nonzero, where $N_{a}$ is the number of strictly positive $\bm{\sigma}$.
@@ IV Algorithmic Approach
Solving Problem 4 in an algorithmically tractable manner can be done through first solving the dual problem using exchange methods to determine the optimal impulse locations and then reconstructing the input through the discretized primal.
@@ IV-A Dual Problem Algorithm
This work leverages an exchange method [21] for solving the semi-infinite program, which consists of iteratively solving a finite version of Problem 10 and adding constraints corresponding to the worst semi-infinite constraint violation. To illustrate this iterative technique, define the modified form of Problem 10, given by
Problem 7
Solve the discretized form of Problem 10 $\mathcal{D}(S)$, where $S\triangleq\{s_{1},s_{2},\dots s_{d}\}\subset\bigcup_{k=1}^{N}T_{k}$ is a finite set, with the addition of the constraint:
$\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{k}\sigma_{k}\leq\Delta V_{tot},$ (15)
where $\Delta V_{tot}=\sum_{k=1}^{N}\Delta V_{k}$ is the total amount of expendable fuel imposed.
Since Problem 10 is reducible according to Lemma 2, then it holds that the finite version of the problem can solve the semi-infinite problem if $S$ is chosen correctly. The additional constraint in Problem 7 prevents artificial infeasibility in the case that the discrete form is under constrained in the exchange method process. This constraint is natural to apply since $\mathcal{D}\leq\mathcal{P}\leq\Delta V_{tot}$ if a feasible point exists, meaning there is no loss in feasibility when this constraint is imposed.
To leverage the reducibility property of the dual, exchange methods gradually add values to the set $S$ such that $\mathcal{D}(S^{(j)})\to\mathcal{D}$ [21, 12]. Several prominent techniques exist in exchange methods for augmenting the discrete set $S$, which effect problem size growth and convergence rate. The technique proposed here simply evaluates the semi-infinite constraint over a grid and adds time points corresponding to the largest constraint violations.
$\begin{gathered}S^{(j+1)}\leftarrow S^{(j)}\cup\operatorname*{arg\,loc\,max}_{\scriptscriptstyle{t\in\cup_{k=1}^{N}T_{k}:f_{k}^{\circ}(\bm{\lambda}_{v}^{(j)}(t))>1+\sigma_{k}^{(j)}}}f_{k}^{\circ}(\bm{\lambda}_{v}^{(j)}(t))\end{gathered}$ (16)
Here, $\bm{\lambda}_{v}^{(j)}(t)=B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{(j)}$ is the primer vector at iteration $j$ of the algorithm. Since $f_{k}^{\circ}(\bm{\lambda}_{v}^{(j)}(t))$ is trivial to compute in this case, this process does not introduce significant computational burden in the algorithmic framework. The discrete set can be initialized with simply the initial and final times, i.e., $S^{(1)}=\{t_{0},t_{f}\}$ or with the optimal times associated with a two-impulse burn as is done in [4]. The selection of the initial times and the addition of new times are hyperparameters of the algorithmic framework which can be adjusted for specific problems, akin to that done in [14].
These algorithmic modifications retain several important features of exchange methods. Namely, it is evident that $\mathcal{D}\leq\mathcal{D}(S^{(j+1)})\leq\mathcal{D}(S^{(j)})$ since $S^{(j)}\subseteq S^{(j+1)}$. Moreover, given a tolerance $\epsilon$, an important result is as follows, which is extended from [4].
@@
Suppose Problem 7 is solved such that $f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})-\sigma_{k}^{*}\leq 1+\epsilon\ \forall t\in T_{k},\forall k\in\mathbb{Z}_{1}^{N}$. Then it follows that $\mathcal{D}(S^{(j)})\leq(1+\epsilon)\mathcal{D}$.
@@
Let $\bar{\bm{\lambda}}_{f}=\bm{\lambda}_{f}^{*}/(1+\epsilon)$ and $\bar{\sigma}_{k}=\sigma_{k}^{*}/(1+\epsilon)$. Evidently, $(\bar{\bm{\lambda}}_{f},\bar{\bm{\sigma}})$ is a feasible point to Problem 10 that satisfies $\mathcal{D}(S^{(j)})/(1+\epsilon)=\bar{\bm{\lambda}}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{k}\bar{\sigma}_{k}\leq\mathcal{D}\implies\mathcal{D}(S^{(j)})\leq(1+\epsilon)\mathcal{D}$. ∎
@@
1: Discretize $\{T_{k}\}_{k=1}^{N}$, $S^{(1)}\leftarrow\{t_{0},t_{f}\}$, $j\leftarrow 1$
2:while $\exists k\in\mathbb{Z}_{1}^{N},t\in T_{k}\ \text{s.t.}\ f^{\circ}_{k}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{(j)})-\sigma_{k}^{(j)}>1+\epsilon$ do
3:  $\bm{\lambda}_{f}^{(j)},\bm{\sigma}^{(j)},\mathcal{D}(S^{(j)})\leftarrowSolve$ Problem 7 with $S^{(j)}$
4:  $S^{(j+1)}\leftarrow S^{(j)}\cup\Delta S$ (Equation 16), $j\leftarrow j+1$
5:end while
6:if $\mathcal{D}(S^{(j)})=\Delta V_{tot}$ then
7:  return Problem Infeasible
8:end if
@@ IV-B Input Reconstruction
Though the above procedure is the most general, it is possible to exploit further algorithmic performance in the case of smooth norm costs and constraints, in which case the direction of the input is known and the ambiguous input magnitudes can be solved through least squares.
Theorem IV.1
Solving the infinite program Problem 4, by means of Algorithm 8 followed by Algorithm 2, is globally convergent to the global minimum to within the imposed tolerance $\epsilon$.
Proof:
Since Problem 7 is convex and bounded above, each subproblem of Algorithm 8 will converge. Moreover, Algorithm 8 will always terminate since $|S^{(j)}|$ is bounded. Finally, since $\mathcal{D}(S^{*})\leq\mathcal{D}(1+\epsilon)$ (Lemma 3), $\mathcal{D}=\mathcal{P}$ (Lemma 1), and $\mathcal{D}(S^{*})=\mathcal{P}(S^{*})$ (strong duality of SOCPs/LPs), $\mathcal{P}(S^{*})\leq\mathcal{P}(1+\epsilon)$. ∎
@@ V Performance and Validation
The problem formulation and corresponding algorithmic approach is validated on the problem of impulsive control for proximity operations of satellites.
@@ V-A System Definition
Relative motion of satellites in orbits subject to single body gravitational force can be modeled as linear time-varying systems under a variety of perturbations, such as drag or spherical harmonics, through a relative orbital element state representation [9, 15], defined by
@@
These states are functions of the classical Keplerian orbital elements of circular two-body motion of the target and chaser spacecraft. The linearized equations of motion under this state are of Jordan form and can be found in [15] along with transformation matrices to RTN frame cartesian state. The ensuing examples use an orbital mean motion of $n=.00113$, consistent with LEO.
@@ V-B Planar Reconfiguration Demonstration
The first example considers a planar rendezvous problem consisting of reducing along track separation from a distance of 100 km to 10 km, i.e., $a\delta\textit{{\oe }}_{0}=[10,100,0,10,0,0]^{\intercal}$ [km] to $a\delta\textit{{\oe }}_{f}=[0,10,0,0,0,0]^{\intercal}$ [km]. The transfer occurs over four orbits, with eight equally spaced time windows enforced over the duration of the transfer with a buffer of $1/10$th of an orbit between each window. Each time window is discretized over 1000 time points in the dual optimization scheme to provide sufficient resolution. The input is scaled such that a unitary impulse corresponds to $1.13$ [m/s] of $\Delta V$ expenditure. Moreover, the norm applied is the $l_{2}$ norm, such that $f_{k}(\cdot)=\|\cdot\|_{2}\ \forall k$.
Figure 1: Planar Reconfiguration Trajectory.
@@
Figure 1 displays the RT-plane relative position, primer vector trajectory, input impulse profile, and dual norm trajectory. Figure 2 shows the trajectory from the perspective of the planar relative orbital elements. Blue dots/arrows denote impulse events, green dots denote starting time, and red dots denote terminal time. The computed dual optimal solution for the problem is $\bm{\lambda}_{f}^{*}=[0.7380,0.0891,-0.2359,-0.7236,0,0]^{\intercal},\bm{\sigma}^{*}=[5.4503,3.0092,3.7714,1.3608,2.0926,0,0.4142,1.5183]^{\intercal}$. The total $\Delta V$ expenditure is $8.860$ [m/s]. Notably, the dual norm exceeds the value of 1 at all but one of the time windows, corresponding exactly to the time windows at which the input constraint is active.
@@
To validate and compare the performance of the proposed algorithm, the problem is also solved under a direct optimization framework whereby the primal problem is discretized with the inputs at each discrete node as solution variables. Under this approach, the number of solution variables increases drastically compared to the dual solution framework, with the total number of solution variables amounting to $n_{u}NM$, with $n_{u}$ being the input dimension, $N$ the number of time windows, and $M$ the discretization count per time window. In contrast, the dual SIP involves $n_{x}+N$ solution variables, with cone count being on the order of $\sim 2N$-$3N$, depending on the problem structure. The poor scalability of direct primal optimization manifests as longer solver times, with this issue being especially apparent for long duration transfers, whereby the discretization induces substantial variable counts.
@@
Under this direct optimization approach, the equivalent problem is solved and produces an equivalent solution. However, the solver runtime is 0.6 seconds, whereas the cumulative solver time for the dual SIP with input reconstruction is 0.1262 seconds [^1] Runtime experiments were performed on an Intel i7‑10750H CPU in MATLAB R2024a with the MOSEK solver [2] and YALMIP modeler [17]. Solver times are determined from extracting the MOSEK solver time from the YALMIP solution. .
V-C VISORS Transfer Demonstration and Comparison